# Luna: An Evaluation Foundation Model to Catch Language Model Hallucinations with High Accuracy and Low Cost

### v1

@@ Abstract
Retriever-Augmented Generation (RAG) systems have become pivotal in enhancing the capabilities of language models by incorporating external knowledge retrieval mechanisms. However, a significant challenge in deploying these systems in industry applications is the detection and mitigation of hallucinations—instances where the model generates information that is not grounded in the retrieved context. Addressing this issue is crucial for ensuring the reliability and accuracy of responses generated by large language models (LLMs) in diverse industry settings. Current hallucination detection techniques fail to deliver accuracy, low latency, and low cost simultaneously. We introduce Luna: a DeBERTA-large (440M) encoder, fine-tuned for hallucination detection in RAG settings. We demonstrate that Luna outperforms GPT-3.5 and commercial evaluation frameworks on the hallucination detection task, with 97% and 96% reduction in cost and latency, respectively. Luna is lightweight and generalizes across multiple industry verticals and out-of-domain data, making it an ideal candidate for industry LLM applications.
@@ 1 Introduction
Large Language Models (LLMs) are broadly used in industry dialogue applications due to their impressive ability to hold a natural conversation and succeed on a variety of reasoning tasks (Zhao et al., 2023). A key challenge in deploying customer-facing LLMs is their propensity for hallucinations, where the model presents cohesive, but factually incorrect information in conversation with a user (Roller et al., 2021; Lin et al., 2022). Retrieval-augmented generation (RAG), a technique for incorporating knowledge relevant to each user query in the LLM prompt, effectively reduces LLM hallucinations in production systems (Lewis et al., 2020). Yet, LLMs still sometimes respond with nonfactual information that contradicts the knowledge supplied by RAG Shuster et al. (2021).
@@
Figure 1: Luna is a lightweight DeBERTA-large encoder, fine-tuned for hallucination detection in RAG settings. Luna outperforms large hallucination detection models (GPT-3.5 and ChainPoll GPT-3.5 ensemble) and RAG evaluation frameworks (RAGAS, Trulens) at a fraction of the cost and millisecond inference speed.
@@
Causes of hallucinations have been extensively studied across different LLM tasks (Zheng et al., 2024; Cao et al., 2022; Das et al., 2022). Key contributing factors include knowledge cutoff (Vu et al., 2023), randomness (Lee et al., 2022), faulty training data (Dziri et al., 2022a; Lin et al., 2022; McKenna et al., 2023), and finetuning with large amounts of new knowledge (Gekhman et al., 2024). Apart from RAG, proposed mitigation solutions explore prompt engineering with chain of thought (Wei et al., 2022), finetuning (Zhang et al., 2024), reinforcement learning with human feedback (Ouyang et al., 2022), and specialized hallucination detection models (Wu et al., 2023; Lin et al., 2022). Gao et al. (2023) and Wu et al. (2023) take it a step further to successfully suppress hallucinations in LLM responses with a detect-and-replace technique. For RAG specifically, evaluation frameworks like RAGAS (Es et al., 2024), Trulens [^1] https://www.trulens.org/, and ARES (Saad-Falcon et al., 2024) have emerged to offer automated hallucination detection at scale. However, these approaches rely on static LLM prompts (RAGAS, Trulens) or finetuning on in-domain data (ARES), which limit their generalization capacity to a breadth of industry applications.
@@
Customer-facing dialogue applications necessitate a hallucination detection system with high-accuracy, low latency, and low cost, such that hallucinations are caught and resolved before reaching the user. Existing LLM prompt-based approaches fail to meet the strict latency requirement due to model size. Moreover, though commericial LLMs like OpenAI’s GPT models (OpenAI, 2023) achieve strong performance, querying customer data through 3rd party APIs is both costly and undesirable for privacy and security reasons. Finetuned BERT-size models can achieve competitive performance to LLM judges (Bohnet et al., 2023; Saad-Falcon et al., 2024; Gao et al., 2023; Li et al., 2024; Yue et al., 2023), offering lower latency and local execution. However, these models require annotated data for finetuning and have not been evaluated for large-scale, cross-domain applications.
@@
Our approach is closest to the concurrently proposed ARES automated RAG evaluation framework (Saad-Falcon et al., 2024), with a few key differences: (1) ARES requires a validation set of in-domain annotated data to finetune a custom evaluation model, while Luna is pre-trained on a broad, cross-domain corpus for built-in generalization; (2) Luna accurately detects hallucinations on long RAG contexts; and (3) Luna is optimized to process up to 16k tokens in milliseconds on deployment hardware.
@@ 3 Luna Model
Figure 2: Distribution of RAG context lengths in our QA RAG training split.
@@ 4.2 Labeling
Table 3: AUROC on the hallucination detection task on the RAG QA test set. Best score in each domain is bolded. Reporting AUROC against GPT-4 annotations. Lunain-domain is our model trained on combined train splits from each domain. LunaOOD is the same model trained on a subset of General Knowledge and Biomed domains.
@@ 7.2 Cost vs Accuracy Trade-offs
API-based hallucination detection methods accrue substantial costs if used continuously in production settings. Luna outperforms GPT-3.5-based approaches while operating at a fraction of the cost. In Figure 1 we illustrate the trade-off between monthly maintenance costs and accuracy for Luna versus our GPT-3.5-based baselines. Costs are estimated assuming average throughput of 10 queries per second, with average query length of 4000 tokens. We use OpenAI API [^5] https://openai.com/api/pricing/ and AWS cloud [^6] https://aws.amazon.com/ec2/pricing/on-demand/ pricing at the time of writing. Detailed cost calculations can be found in Appendix A.
@@ 8.2 Future Work
Appendix A Cost Calculations
Costs are estimated assuming average throughput of 10 queries per second (qps), with average query length of 4000 tokens, and NVIDIA L4 GPU deployment hardware.
@@ Appendix A Cost Calculations
At the time of writing, querying GPT-3.5-turbo through OpenAI API costs \$0.50 / 1M input tokens and \$1.50 / 1M output tokens [^8] https://openai.com/api/pricing/. In our test set, we observe the average output token length from GPT-3.5 at 200 tokens. Using average input length of 4000 tokens, the cost of a single query is
@@
Wuth ChainPoll ensemble, we request 3 outputs per query, bringing the cost of a single query up to
@@
Appendix B Latency Optimizations

### v2

@@ Abstract
Retriever-Augmented Generation (RAG) systems have become pivotal in enhancing the capabilities of language models by incorporating external knowledge retrieval mechanisms. However, a significant challenge in deploying these systems in industry applications is the detection and mitigation of hallucinations—instances where the model generates information that is not grounded in the retrieved context. Addressing this issue is crucial for ensuring the reliability and accuracy of responses generated by large language models (LLMs) in diverse industry settings. Current hallucination detection techniques fail to deliver accuracy, low latency, and low cost simultaneously. We introduce Luna: a DeBERTA-large (440M) encoder, fine-tuned for hallucination detection in RAG settings. We demonstrate that Luna outperforms GPT-3.5 and commercial evaluation frameworks on the hallucination detection task, with 97% and 91% reduction in cost and latency, respectively. Luna is lightweight and generalizes across multiple industry verticals and out-of-domain data, making it an ideal candidate for industry LLM applications.
@@ 1 Introduction
Large Language Models (LLMs) are broadly used in industry dialogue applications due to their impressive ability to hold a natural conversation and succeed on a variety of reasoning tasks (Zhao et al., 2023). A key challenge in deploying customer-facing LLMs is their propensity for hallucinations, where the model presents cohesive, but factually incorrect information in conversation with a user (Roller et al., 2021; Lin et al., 2022). Retrieval-augmented generation (RAG), a technique for incorporating knowledge relevant to each user query in the LLM prompt, effectively reduces LLM hallucinations in production systems (Lewis et al., 2020). Yet, LLMs still often respond with nonfactual information that contradicts the knowledge supplied by RAG Shuster et al. (2021); Magesh et al. (2024).
@@
Figure 1: Luna is a lightweight DeBERTA-large encoder, fine-tuned for hallucination detection in RAG settings. Luna outperforms zero-shot hallucination detection models (GPT-3.5, ChainPoll GPT-3.5 ensemble) and RAG evaluation frameworks (RAGAS, Trulens) at a fraction of the cost and millisecond inference speed.
@@
Causes of hallucinations have been extensively studied across different LLM tasks (Zheng et al., 2024; Cao et al., 2022; Das et al., 2022). Key contributing factors include knowledge cutoff (Vu et al., 2023), randomness (Lee et al., 2022), faulty training data (Dziri et al., 2022a; Lin et al., 2022; McKenna et al., 2023), and finetuning with large amounts of new knowledge (Gekhman et al., 2024). Apart from RAG, proposed mitigation solutions explore prompt engineering with chain of thought (Wei et al., 2022), finetuning (Zhang et al., 2024), reinforcement learning with human feedback (Ouyang et al., 2022), and specialized hallucination detection models (Wu et al., 2023; Lin et al., 2022). For RAG specifically, evaluation frameworks like RAGAS (Es et al., 2024), Trulens [^1] https://www.trulens.org/, and ARES (Saad-Falcon et al., 2024) have emerged to offer automated hallucination detection at scale. However, these approaches rely on static prompts (RAGAS, Trulens) or finetuning on in-domain data (ARES), which limit their capacity to generalize to a breadth of industry applications. Gao et al. (2023) and Wu et al. (2023) take it a step further to successfully suppress hallucinations in LLM responses with a detect-and-replace technique. Though, due to prohibitively slow latency of their LLM evaluation models, real-time hallucination prevention in production systems still remains a challenge.
@@
Customer-facing dialogue applications necessitate a hallucination detection system with high-accuracy, low cost, and low latency, such that hallucinations are caught and resolved before reaching the user. Few/zero-shot LLM approaches fail to meet the strict latency requirement due to model size. Moreover, though commericial LLMs like OpenAI’s GPT models (OpenAI, 2023) achieve strong performance, querying customer data through 3rd party APIs is both costly and undesirable for privacy and security reasons. Finetuned BERT-size models can achieve competitive performance to LLM judges (Bohnet et al., 2023; Saad-Falcon et al., 2024; Gao et al., 2023; Li et al., 2024; Yue et al., 2023), offering lower latency and local execution. However, these models require annotated data for finetuning and have not been evaluated for large-scale, cross-domain applications.
@@
Our approach is closest to the concurrently proposed ARES automated RAG evaluation framework (Saad-Falcon et al., 2024), with a few key differences: (1) ARES requires a validation set of in-domain annotated data to finetune a custom evaluation model, while Luna is pre-trained on a cross-domain corpus for built-in generalization; (2) Luna accurately detects hallucinations on long RAG contexts; and (3) Luna is optimized to process up to 16k tokens in milliseconds on deployment hardware.
@@ 3 Luna Model
Figure 2: Distribution of RAG context token lengths in our QA RAG training split.
@@ 4.2 Labeling
Table 3: AUROC on the hallucination detection task on the RAG QA test set. Best score in each domain is bolded. Lunain-domain is our model trained on combined train splits from each domain. LunaOOD is the same model trained on a subset of General Knowledge and Biomed domains.
@@ 7.2 Cost vs Accuracy Trade-offs
API-based hallucination detection methods accrue substantial costs if used continuously in production settings. Luna outperforms GPT-3.5-based approaches while operating at a fraction of the cost. In Figure 1 we illustrate the trade-off between monthly maintenance costs and accuracy for Luna versus our GPT-3.5-based baselines. Costs are estimated assuming average throughput of 10 queries per second, with average query length of 4000 tokens. We use OpenAI API [^5] https://openai.com/api/pricing/ and AWS cloud [^6] https://aws.amazon.com/ec2/pricing/on-demand/ pricing at the time of writing. Detailed cost calculations can be found in Appendix B.
@@ 8.2 Future Work
Appendix A Response Generation Prompt
We use the following prompt template to generate LLM responses for each sample in our QA RAG dataset. Context documents, separated by line breaks, along with the question are slotted in for each generation sample.
Use the following pieces of context to answer the question.
{documents}
Question: {question}
Appendix B Cost Calculations
Costs are estimated assuming average throughput of 10 queries per second (qps), with average RAG query length of 4000 tokens, and NVIDIA L4 GPU deployment hardware. When estimating LLM cost for >1qps we assume concurrency is implemented to process multiple queries in parallel.
@@ Appendix B Cost Calculations
At the time of writing, querying GPT-3.5-turbo through OpenAI API costs \$0.50 / 1M input tokens and \$1.50 / 1M output tokens [^8] https://openai.com/api/pricing/. In our test set, we observe the average output token length from GPT-3.5 at 200 tokens. Using average input length of 4000 tokens, the cost of a single query is roughly
@@
With ChainPoll ensemble, we request 3 outputs per query, bringing the cost of a single query up to
@@
RAGAS Costs
RAGAS makes 2 OpenAI API calls per an input RAG example. The first query extracts a list of claims from the response. The second requests the LLM to evaluate the faithfulness of each extracted claim to the RAG context. We estimate that the output length of the first query is roughly equal to the length of the RAG response; and the output length of the second query is roughly 3x the length of the response, since it includes the original claims followed by a faithfulness score and an explanation. Factoring in overhead token length of each prompt, we calculate the cost per query to be
$Query1=\$380/1M$ (11)
$Query2=\$2730/1M$ (12)
Then, the monthly cost of serving 10qps is:
$93710qps*2,592,000*(\$380+\$2730)/1M=\$79,$ 937 (13)
Trulens Costs
Trulens makes 1 OpenAI per each sentence in the response. For this calculation, we estimate 3 sentences per response, which aligns with our obesrvations on the QA RAG dataset. Each query returns original sentence, a groundedness score (1-10), and an explanation. Here we assume that the token length of the explanation is roughly equal to the token length of the input sentence. The cost of a single query is roughly
$(4k*\$0.5+2*75*\$1.5)/1M=\$0.0022$ (14)
Using 2,592,000 seconds/month, the monthly cost of serving 10qps with Trulens is:
$01610qps*2,592,000*3*\$0.0022=\$173,$ 016 (15)
Appendix C Latency Optimizations
@@
Appendix D Latency Comparison
We empirically estimate the latency of Luna and each baseline model. Luna latency is discussed in Appendix C. For LLm models that query OpenAI API, we calculate the average latency per query after querying the API multiple times with an input of 4000k tokens, split between 3800 tokens for the context, 25 tokens for the question, and 75 tokens for the response.
Model s/4k %change
Luna 0.23 -
GPT-3.5 2.5 -91%
ChainPoll n=3 3.0 -93%
Trulens 3.4 -93%
RAGAS 5.4 -96%
Table 6: Model latency (in seconds), comparing Luna to LLM baselines. We also report the % difference between Luna and LLM-based models.